OpenStudy (anonymous):

Find an equation in standard form for the hyperbola with vertices at (0, ±10) and asymptotes at y = ±5/4 x.

OpenStudy (anonymous):

center is at the origin, half way between the vertices, so you know it looks like \[\frac{x^2}{a^2}-\frac{y^2}{b^2}=1\]

OpenStudy (anonymous):

\(a=10\implies a^2=100\) so now we have \[\frac{x^2}{100}-\frac{y^2}{b^2}=1\]

OpenStudy (anonymous):

for \(b\) note that \(\frac{a}{b}=\frac{10}{b}=\frac{5}{4}\) and solve for \(b\)

OpenStudy (anonymous):

b=12, y^2/100-x^2/144? @satellite73

OpenStudy (jhannybean):

\[\large \frac{10}{b}=\frac{5}{4}\] cross multiply..

OpenStudy (anonymous):

okay, wrong number..opps. b5=14.. 11?did i do that right?

OpenStudy (anonymous):

@Jhannybean

OpenStudy (jhannybean):

what is 4 times 10?

OpenStudy (jhannybean):

\[\large 10 + 4 \ne 10 \cdot 4\]

OpenStudy (anonymous):

40..

OpenStudy (jhannybean):

and what is 40 divided by 5?

OpenStudy (anonymous):

8

OpenStudy (anonymous):

we're getting there right?

OpenStudy (anonymous):

yeah, It's been a really long day..haha

OpenStudy (anonymous):

so a is 5 and b is 8?

OpenStudy (anonymous):

i can tell you should solve \[\frac{10}{b}=\frac{5}{4}\] in your head by noting that \(10=2\times 5\) and so \(b=2\times 4\)

OpenStudy (anonymous):

i.e. \(b=8\) and so the equation is \[\frac{x^2}{100}-\frac{y^2}{64}=1\] lets check it

OpenStudy (anonymous):

oops i had it backwards !!

OpenStudy (anonymous):

\[\frac{y^2}{100}-\frac{x^2}{64}=1\]

OpenStudy (anonymous):

that's better http://www.wolframalpha.com/input/?i=hyperbola+y^2%2F100-x^2%2F64%3D1

OpenStudy (anonymous):

okay I understand now thanks guys!